From past experience, a professor knows that the test score of a student taking her final examination is a random variable with mean 70 and variance 8.  What is the upper bound of the probability that a student will score between 65 and 75


1
Expert's answer
2022-01-12T16:38:32-0500

Chebyshev’s inequality states that for any probability distribution of an rv "X" and any number "k" that is at least "1,P(|X-\mu|\geq k\sigma)\leq 1/k^2."

"P(65\leq X\leq75)=P(|X-70|\leq 5)"

"=1-P(|X-70|>5)\geq 1-\dfrac{\sigma^2}{5}"

"=1-\dfrac{8}{5^2}=0.68"